OpenStudy (anonymous):

What are the x-intercept(s) of the graph of the equation y − x2 = −4x − 5?

OpenStudy (anonymous):

y=0 in x-intercept \[y-x^{2}=-4x-5\] \[y=x^{2}-4x-5\] y=0 in x intercept. \[0=x^{2}-4x-5\] \[0=x^{2}+x-5x-5\] \[0=x(x+1)-5(x+1)\] \[0=(x-5)(x+1)\] so, x=5 or x=-1